Compute for the mean of the set of numbers below. Write your answer on the

space provided in each item. Round off your answers to two decimal places.

1) 2, 5, 7, 3, 9

2) 12, 15, 7, 9, 10, 17

3) 3, 5, 7, 15, 17, 19, 21

4) 20, 32, 26, 22, 27, 30, 28, 21

5) 30, 35, 29, 34, 37, 40, 43, 36, 38



1
Expert's answer
2021-05-07T09:38:20-0400

1) 2, 5, 7, 3, 9


"\text{Mean} =\dfrac{\text{sum of observation}}{\text{Total Observation}}=\dfrac{2+5+7+3+3+9}{5}=\dfrac{26}{5}=5.2"



2) 12, 15, 7, 9, 10, 17


"\text{Mean} =\dfrac{\text{sum of observation}}{\text{Total Observation}}=\dfrac{12+15+7+9+10+17}{6}=\dfrac{70}{6}=11.66"


3) 3, 5, 7, 15, 17, 19, 21


"\text{Mean} =\dfrac{\text{sum of observation}}{\text{Total Observation}}=\dfrac{3+5+7+15+17+19+21}{7}=\dfrac{87}{7}=12.42"


4) 20, 32, 26, 22, 27, 30, 28, 21


"\text{Mean} =\dfrac{\text{sum of observation}}{\text{Total Observation}}=\frac{20+32+26+22+27+30+28+21}{8}=\dfrac{206}{8}=25.75"



5) 30, 35, 29, 34, 37, 40, 43, 36, 38


"\text{Mean} =\dfrac{\text{sum of bservation}}{\text{Total Observation}}=\frac{30+35+29+34+37+40+43+36+38}{9}=\dfrac{322}{9}=35.77"